OpenStudy (anonymous):

Find the value of k so that the graph of 8y +kx=4 and the line containing the point (5,-8) ans (2,4) are parallel

OpenStudy (aum):

What is the slope of the line through the given two points?

OpenStudy (anonymous):

i didnt get one

OpenStudy (aum):

Slope of the line through two points = \(\Large\frac{y_2-y_1}{x_2-x_1} \)

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

the other is -8

OpenStudy (anonymous):

??

OpenStudy (aum):

Correct. What is the slope of the line 8y +kx=4 in terms of k? Put it in the slope-intercept form first (that is, y = mx + b)

OpenStudy (aum):

The -4 is correct for the slope of the line through the given two points.

OpenStudy (anonymous):

not sure what to do

OpenStudy (aum):

8y + kx = 4 Yo find the slope of the above line, put it in slope-intercept form: subtract kx 8y = -kx + 4 divide by 8 y = -k/8 * x + 4/8 The coefficient of the x-term is the slope. slope = -k/8 Slope of the line through (5,-8) and (2,4) is -4 Since the two lines are parallel, their slopes must be the same. -k/8 = -4 -k = -32 k = 32

OpenStudy (anonymous):

oooohhhh i usually do -a/b from standard form to get slope i put -1 instead of -k

OpenStudy (aum):

Yes, if you know how to calculate slope directly from the standard form you can skip that step. But most students are familiar with finding slope from the slope-intercept form.

OpenStudy (aum):

Here a = k and b = 8 slope = -a/b = -k/8 (same as what I got above)

OpenStudy (aum):

a is the coefficient of x and b is the coefficient of y in the standard form.

OpenStudy (anonymous):

yup i just saw that, the x and y where switched

OpenStudy (aum):

yeah, got to watch out for such things.

OpenStudy (anonymous):

yea do we ever need to find c

OpenStudy (anonymous):

i haven't seen a problem like that

OpenStudy (aum):

by "c" do you mean the y-intercept?

OpenStudy (anonymous):

Ax+bx=c

OpenStudy (aum):

8y +kx=4. Here c is already known. It is 4. So not needed here.

OpenStudy (anonymous):

ok i get it, here another problem easy but i dont get the concept since i get another solution

OpenStudy (anonymous):

one maid clean the house in 6 hours. another maid can do the job in 5 hours. how long will it take them to do the working job together.

OpenStudy (anonymous):

30/11

OpenStudy (aum):

correct.

OpenStudy (anonymous):

but i thought it was like this 1/5+1/6

OpenStudy (aum):

Both are true. In 1 hour, the first maid can clean 1/6th of the house. In 1 hour, the second maid can clean 1/5th of the house. If they work together, then in 1 hour they can clean (1/6 + 1/5)th of the house Or (11/30)th of the house. So to clean the whole house it will take 30/11 hours.

OpenStudy (aum):

11/30 th of the house takes 1 hour 1 house will take 1 * 1 divided by 11/30 or 30/11 hours.

OpenStudy (aum):

If 1/4th of the house can be cleaned in 1 hour, then the house will take 4 hours. If 1/13th of the house can be cleaned in 1 hour, then the house will take 13 hours. If 1/x th of the house can be cleaned in 1 hour, then the house will take x hours. If p/q th of the house can be cleaned in 1 hour, then the house will take q/p hours. If 11/30 th of the house can be cleaned in 1 hour, then the house will take 30/11 hours.

OpenStudy (anonymous):

ok that make sense

OpenStudy (aum):

alright.

OpenStudy (anonymous):

thank you

OpenStudy (aum):

you are welcome.